OpenStudy (anonymous):

Is this a Exponential Growth or Exponential Decay, y = 0.65(1.3)^(x)

ganeshie8 (ganeshie8):

exponent positive or negative ?

OpenStudy (anonymous):

its just x so im guessing positive

ganeshie8 (ganeshie8):

as you increase x, what happens to y ?

OpenStudy (anonymous):

it should decrease rite

OpenStudy (anonymous):

depends entirely on whether the base (in this case 1.3) is greater than or less than one

OpenStudy (anonymous):

since \(1.3>1\) this models growth

OpenStudy (anonymous):

so is y=1 or x=1

ganeshie8 (ganeshie8):

@satellite73 put x = 1 & x = -1 , compare the values, and tell me if it you still say it doesnt depend on exponent sign :)

OpenStudy (anonymous):

?

OpenStudy (unklerhaukus):

\[ y(x) = 0.65(1.3)^{x}\] \[ y(1) = 0.65(1.3)^{1}=0.845\]\[ y(10) = 0.65(1.3)^{10}\approx8.96\] as x has grown, y has grown

OpenStudy (anonymous):

you lost me entirely \(y=P_0\times b^x\) grows so long as \(b>1\) and decreases if \(b<1\)

OpenStudy (anonymous):

in other word, in this example, if \[x_1<x_2\implies .65(1.3)^{x_1}<.65(1.3)^{x_2}\]

OpenStudy (anonymous):

here is a nice picture showing it going up http://www.wolframalpha.com/input/?i=y+%3D+0.65%281.3%29^%28x%29